OpenStudy (1645323):

The U.S. Postal Service has hired you to construct a box (top included) such as the one below that minimizes surface area to reduce cost of materials. The box must have a square base and the volume of the box must be 5,000 inches3. What is the smallest surface area this box can have, to the nearest whole number? Select one: a. 1,412inches2 b. 1,754 inches2 c. 1,686 inches2 d. 1,238 inches2

OpenStudy (1645323):

@ganeshie8 @thomaster @AccessDenied @beccaboo333

OpenStudy (1645323):

@SolomonZelman

OpenStudy (solomonzelman):

rectangular box. The volume is 5000"³ and what is the smallest surface area. (just rephrasing for myself. ) Just 2 examples of dimensions. 1) lets take 20, 25, 10, the volume = 20×25×10=5000 Surface area = 2(20×25)+2(20×10)+2(10×25)=1,900 2) NOW, lets take 5000, 1, 1 the volume = 5000×1×1 Surface area = 2(5000×1)+2(1×1)+2(1×1)=10,004 See? the smaller the dimensions overall are, the smaller the surface

OpenStudy (1645323):

@SolomonZelman is 10,004/6 the answer?

OpenStudy (solomonzelman):

10,004/6 ? How do you get that ?

OpenStudy (1645323):

the surface area and divide it by 6 for each face

OpenStudy (solomonzelman):

the surface area (of a rectangular box) with what dimensions ?

OpenStudy (1645323):

xo non e of the answer choices are correct?

OpenStudy (solomonzelman):

it is asking you, what is the smallest surface area that a 5000 box can have ?

OpenStudy (solomonzelman):

the smallest surface area, would be the one that has the closest-to-each-other (3 numbers) the product of which gives 5000.